3a. In last week’s football game, Ralph scored 6 less than twice what Tommy scored. The sum of their scores is 30. How many points did Ralph and Tommy score individually?

3b. The ages of three sisters are three even consecutive integers. If the sum of the 1st, four times the 2nd, and twice the 3rd is 86, what are the three ages?

3c. You are buying a computer at a markdown price of $960. If the markdown price was 20% off the original price, how much was the computer originally?
1
Expert's answer
2013-01-23T10:57:14-0500
3a. In last week's football game, Ralph scored 6 lessthan twice what Tommy scored. The sum of their scores is 30. How many points
did Ralph and Tommy score individually?
Answer.
Let R be the scores of Ralph, and T be the scores ofTommy.
Then
R = 2T-6
and
R+T = 30
Therefore
30 = R+T =2T-6+T = 3T-6
3T = 36
T=12
whence
R = 30-T = 18
Thus Ralph scores 18, and Tommy 12
3b. The ages of three sisters are three even consecutiveintegers. If the sum of the 1st, four times the 2nd, and twice the 3rd is 86,
what are the three ages?
Answer.
Let X be the age of the youngest sister.
Then two other sisters have ages X+2 and X+4 respectively.
We also know that
X + 4(X+2) + 2(X+4) = 86
X + 4X + 8 + 2X + 8 = 86
7X = 86 - 16 = 70
X = 10
So the ages of sisters are
10, 12, 14
3c. You are buying a computer at a markdown price of$960. If the markdown price was 20% off the original price, how much was the
computer originally?
Answer.
Let X be the original price.
Then the markdown price constitute 80% of X, that is
960 = 0.8 X
whence
X = 960/0.8 = $1200